OpenStudy (anonymous):

Let f(x) = x-2/4 and g(x)= 2x^2+4. I already found f(g(x)). Now I need help finding g(f(x)).

OpenStudy (anonymous):

f(g(x)) = f(2x^2 + 4) = (2x^2 + 4 - 2)/4 = (2x^2 + 2)/4 = 2(x^2 + 1)/4 =(x^2 + 1)/2 .........now I need g(f(x)).

OpenStudy (anonymous):

so far I have gotten g(f(x))= g((x-2)/4)= 2x^2+4(-2/4)= 2x^2+(-8/16)= and I am stuck

OpenStudy (anonymous):

@mathstudent55 I opened a new one because the other got confusing and long.

OpenStudy (anonymous):

\[f(x)=\frac{x-2}{4}\\ g(x)=2x^2+4\\ \begin{align*}g(f(x))&=g\left(\frac{x-2}{4}\right)\\ &=2\left(\frac{x-2}{4}\right)^2+4\\ &=2\left(\frac{(x-2)^2}{16}\right)+4\\ &=\frac{1}{8}\left(x^2-4x+4\right)+4\\ &=\frac{1}{8}x^2-\frac{1}{2}x+\frac{1}{2}+4\\ &=\frac{1}{8}x^2-\frac{1}{2}x+\frac{9}{2} \end{align*}\]